myininaya (myininaya):

Is it me or all of us? I can't highlight anything on the site, you know to copy it and paste it elsewhere.

thomaster (thomaster):

everywhere? or does it only occur on openstudy? Never had this problem

OpenStudy (charlotte123):

Don't have this problem at all! Can copy and paste just fine ^_^

myininaya (myininaya):

Just this site.

myininaya (myininaya):

So you copy what I just said and paste it somewhere else?

myininaya (myininaya):

I can copy this from Wikipedia: In mathematics, the factorial of a non-negative integer n, denoted by n!, is the product of all positive integers less than or equal to n. For example, 5! = 5 \times 4 \times 3 \times 2 \times 1 = 120. \ The value of 0! is 1, according to the convention for an empty product.[1] The factorial operation is encountered in many areas of mathematics, notably in combinatorics, algebra and mathematical analysis. Its most basic occurrence is the fact that there are n! ways to arrange n distinct objects into a sequence (i.e., permutations of the set of objects). This fact was known at least as early as the 12th century, to Indian scholars.[2] The notation n! was introduced by Christian Kramp in 1808.[3] The definition of the factorial function can also be extended to non-integer arguments, while retaining its most important properties; this involves more advanced mathematics, notably techniques from mathematical analysis. And paste it here as I did. But I can't collect this text here and paste it somewhere else.

myininaya (myininaya):

I get no blue And it is sad news

thomaster (thomaster):

try closing the browser and start it up again.

myininaya (myininaya):

thomaster Best Response Medals 0 everywhere? or does it only occur on openstudy? Never had this problem omg you are a wizard

thomaster (thomaster):

haha no, rebooting something usually fixes a lot of problems :P

myininaya (myininaya):

It was weird because I didn't have any trouble copying text in other places.

myininaya (myininaya):

Thanks for the help you guys.